7x=6y+24 I need to find the slope, can someone please help me!!
Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)


Hi,
*Note: The standard point-slope form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
Solving for y to put into the slope-intercept format to determine the slope
7x=6y+24
7x - 24 = 6y
(7/6)x - 4 = y
m = (7/6) Note: when y increases by 7, then x increases by 6


Answer by Alan3354(69443)

This type of problem is made to seem more complex than it is.
7x=6y+24
"Put on slope-intercept form" means solve for y.
7x=6y+24
6y = 7x - 24
y = (7/6)x - 4
The slope, m, is the coefficent of x
m = 7/6
That's all it takes.
-------------
PS the y-intercept is the -4, or the point (0,-4)